OpenStudy (anonymous):

F(x)=x^3 + x + 4 What is g'(6)?

sam (.sam.):

f'(6) or g'(6)

OpenStudy (anonymous):

Sorry, I forgot to mention that g(x) is the inverse of f(x)

OpenStudy (anonymous):

So yes, g'(6)

OpenStudy (amistre64):

1/f'(x) tends to be defined as the inverse

OpenStudy (amistre64):

so, whats our f'(x)?

OpenStudy (anonymous):

3x^2 +1

OpenStudy (anonymous):

yup thanks for that little piece of info, I just forgot that part, I think I might be able to solve it ^_^

OpenStudy (amistre64):

and x=6 thats what; 36(3) + 1 = 90+18+1

OpenStudy (amistre64):

yep, an inverse flips things about the y=x line so a straight line of a slope just flips as well

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

I'm looking for g'(6)

OpenStudy (amistre64):

right; and since g'(x) is the inverse of f'(x); g'(x) = 1/f'(x)

OpenStudy (anonymous):

so far I have 1/(3x^2 + 1) = g(x)

OpenStudy (anonymous):

so I find g'(x) from that and plugin 6?

OpenStudy (amistre64):



OpenStudy (amistre64):

you find g'(6) by finding 1/f'(6)

OpenStudy (amistre64):

g'(6) = 1/f'(6) if you know f'(6) then reciprocal it

OpenStudy (anonymous):

ohhhh sorry I thought 1/f'(x) = g(x)

OpenStudy (anonymous):

1/109?

OpenStudy (amistre64):

think about it like this; a slope (y/x) IS a vector: <x,y> when we flip things over the y=x line we get: <y,x> as the new slope

OpenStudy (anonymous):

lol mind if we do this algebraically w/o a calculator? I'm taking a practice exam from section 1 so it would be cool if I could learn it like that

OpenStudy (amistre64):

3(36) + 1 = 90 + 18 + 1 = 90+10+9 = 109 so yes, g'(6) = 1/109

OpenStudy (anonymous):

Its weird, that's not one of the multiple choice answers, did I type something wrong?

OpenStudy (amistre64):

maybe

OpenStudy (anonymous):

Can you view the png attachment?

OpenStudy (amistre64):

x^3 -> 3x^2 x -> 1 4 -> 0 f'(x) = 3x^2 + 1 the derivative is right; and the original f(x) doesnt undo very well to check its inverse with which is why this method is perfered

OpenStudy (amistre64):

the inverse of a slope is simply swap top to bottom: y/x inverts to x/y

OpenStudy (amistre64):

3(36) = 108 + 1 = 109 the flip of that is 1/109 let me see if im forgetting something with this tho ....

OpenStudy (anonymous):

yea lol i found this on paul's notes, maybe it will be helpful? http://tutorial.math.lamar.edu/Classes/CalcI/DiffInvTrigFcns.aspx

OpenStudy (amistre64):

yeah, lets do a problem I can test :) y = x^2 ; y = sqrt(x) y' = 2x ; y' = 1/2sqrt(x) say x=4 y' = 8 ; y' = 1/4 hmmm, yeah, im missing something

OpenStudy (amistre64):

f'(f(x)) eh 1/2(sqrt(x)) would be f'(f(x)) in this case

OpenStudy (anonymous):

sqrt is square root right?

OpenStudy (amistre64):

yeah

OpenStudy (amistre64):

given that little side track :/ y = x^3+x+4 y-4 = x^3+x doesnt invert that well

OpenStudy (amistre64):

http://oregonstate.edu/instruct/mth251/cq/Stage6/Lesson/inverseDeriv.html heres what I was recalling tho

OpenStudy (anonymous):

lol this is so confusing

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=invert+x%5E3%2Bx%2B4 now THAT would be a mess

OpenStudy (amistre64):

im forgetting something pretty simple in all this ...

OpenStudy (amistre64):

what is f(6)? by chance?

OpenStudy (anonymous):

226

OpenStudy (amistre64):

216+6+4 = 226 (6,226) the inverted point would then be (226,6)

OpenStudy (anonymous):

so g(226) = f(6) ?

OpenStudy (amistre64):

maybe y=x^2 ; x=4, y=16 (4,16) inverts to (16,4) 2(4) = 8 ; but the inverse 1/2(sqrt(x)) needs to be 1/8 to make sense of this ugh .... it bites getting old lol

OpenStudy (anonymous):

hmm i have an idea

OpenStudy (anonymous):

shouldnt we set the equation = to 6?

OpenStudy (anonymous):

Solve for x Then find the derivative of the function and use it to finish solving for g'(6)

OpenStudy (amistre64):

\[f(f^{-1}(x))=x\] \[f(y)=x\] \[f'(y)\ y'=1\] \[y'=\frac{1}{f'(y)}\] \[y'=\frac{1}{f'(f(x))}\]

OpenStudy (amistre64):

yes, f(6) = 226 3(226)^2 + 1 tho is not going to get us one of those answers is it

OpenStudy (anonymous):

Hmm...

OpenStudy (amistre64):

y = f(x) f(y) = x ; is our inverse still comes out the same in the end doesnt it

OpenStudy (anonymous):

comes out in the same end?

OpenStudy (amistre64):

the inverse of a slope is its reciprocal; plain and simple. if the slope at f(6) = 209; then the slope of its inverse is 1/209 .... but either this is thinking askew or the problem you have is in error

OpenStudy (amistre64):

x^2 + x + 4 2x + 1 13; 1/13 that aint it -x^3 + x + 4 -3x^2 + 1 -107 ... not it arghh!!

OpenStudy (anonymous):

Hmm. These questions are annoying, I wish I had an answer key. Hey well, thanks anyway! I'll move on to other practice problems and maybe I'll ask my teacher the next day.

OpenStudy (amistre64):

ok, good luck with it all :)